# PersuasiveToM: A Benchmark for Evaluating Machine Theory of Mind in Persuasive Dialogues

### v1

@@ 4.1 Baseline Models
We evaluate PersuasiveToM on eight frontier LLMs from different sources and with different levels of capabilities: Llama-3.1-8B-Instruct Dubey et al. (2024), Qwen-2.5-7B-Chat Yang et al. (2024), Gemma-2-9B-it Team et al. (2024), GLM4-9B -ChatGLM et al. (2024), Mixtral-8x7b-Instruct Jiang et al. (2024), and ChatGPT-series(GPT-4o-mini, GPT-4o-0806). By following the common practices (Kim et al. (2023); Sabour et al. (2024)), we test these models with two types of prompts: (1) vanilla zero-shot prompting directly asks LLMs to give a choice without any explanation; (2) CoT prompting method by following Kojima et al. (2022) and using the prompt “Let’s think step by step.” to elicit the reasoning process and extract the choices by string matching. The temperature for generating answers is set to 0.7111LLMs occasionally output with illegal format. We choose a low but nonzero temperature to resample the answers for these invalid generations.. To measure the specific performance gap between humans and the state-of-the-art machine on the PersuasiveToM, we employ three graduate students in computer science to complete the human evaluation task. To avoid the bias of LLMs toward a specific choice letter, we shuffle the choices to maintain a nearly uniform distribution of correct choices over the dataset. Prompts used for vanilla zero-shot prompting and CoT prompting are shown in Appendix B.2.
@@ 5.1 Main Results
Our results show that smaller models, such as Gemma-2-9B and Qwen-2.5-7B, can perform reasonably well in inferring the persuader’s desires, achieving an accuracy of over 98%, which is competitive with GPT-4o. This suggests that most LLMs can easily discern the desires of the persuader. However, when it comes to the desires of the persuadee, performance is relatively lower. Unlike the static desires of the persuader, the persuadee’s desires are dynamic, evolving from an initial state to a final state, often with neutral expressions in between. This lower performance highlights that inferring the dynamic desires of the persuadee remains a significant challenge for LLMs. In summary, LLMs are better at inferring the desires of the persuader than those of the persuadee.
@@
On belief questions, larger models like GPT-4o perform much better than smaller models on reasoning the beliefs of both parties. The performance difference between the reasoning persuader’s beliefs and persuadee’s beliefs is subtle. This is because both parties’ beliefs dynamically change with each other’s speech during the conversation, unlike the desire of the persuader which has a more obvious and precise tendency. The difficulty of reasoning persuader and persuadee’s belief is similar.
@@
Strategy Prediction and judgement.
@@
Both ToM reasoning and ToM application tasks indicate that CoT reasoning has not consistently improved performance, yet improve strategy prediction to some extent for most LLMs. We attribute this limited improvement to the inherent lack of ToM capability in LLMs. These models struggle to simulate the human cognitive reasoning process, which affects their ability to generate correct answers in tasks requiring a nuanced understanding of the mental state.
@@
To obtain a baseline for human performance, we recruited participants to complete the questions. More details of human evaluation are shown in Appendix B.1. As shown in Table 4, our human participants outperformed LLMs on all tasks. In particular, although GPT-4 reaches close performance in humans, it still falls short of understanding and reasoning the complex dynamics such as the intention of persuaders and the desire of persuadee, which involves complex psychological changes, highlighting a significant gap in current LLMs and humans.
@@ 5.2 In-depth Analysis
Figure 3 summarizes the distribution of errors of Desire questions happening in different stages of dialogue progress with and without CoT reasoning. The error distribution for persuader and persuadee has a significant difference. At the beginning of the dialogue, LLMs may not accurately understand the persuader’s desire, but as the dialogue progresses, the persuader’s desire becomes relatively easy to identify. However, for the persuadee, the desire to reject at the early stage of the dialogue is relatively easy to recognize. As the persuasion proceeds, the persuadee may begin to contemplate and hesitate over the persuader’s proposal, leading to complex and nuanced psychological activities that make it difficult for the LLM to accurately judge the persuadee’s desire. As the dialogue approaches its end, the persuadee shows a tendency to agree, making the reasoning of desire easier. This suggests LLMs still fall short in ToM reasoning regarding desire shifts.
@@
Figure 4 and 5 summarize error types of belief questions for each model with and without CoT. We use Distil-BERT [^2] https://huggingface.co/distilbert/distilbert-base-uncased-finetuned-sst-2-english Sanh (2019) to discriminate whether the choice of LLMs has the same attitude polarity to the ground-truth. Interestingly, we found that the proportion of errors with the same polarity is higher when predicting the persuader’s beliefs, whereas the opposite trend was observed when predicting the persuadee’s beliefs. This highlights reasoning shifting mental states is still challenging for LLMs.
@@
Given the high error rate in intention questions related to the persuader, we conducted an analysis of the error types. Our findings reveal that most LLMs exhibit a bias toward predicting intentions characterized by making the other person feel accepted through concessions, promises, or benefits. Figure 6 illustrates the proportion of errors resulting from misclassifying intentions into this category. We hypothesize that this bias may stem from the pretraining phase, particularly with Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2017), which tends to prioritize safety and politeness. This may explain the models’ bias toward predicting intentions emphasizing benefits and concessions, even when misaligned with the dialogue context. We provide a case study in Appendix B.3.
5.3 How Well LLMs Track Mental States of Persuadees?
To study the ability of LLMs to track mental states and completely understand the psychological dynamics in the dialogue. We evaluate the consistency of LLMs. To assess consistency, we measure whether the model maintains a stable understanding of the persuadee’s beliefs, desires, and intentions across multiple turns in the conversation. Specifically, we indicate a success only when all the questions in the dialogue are answered correctly.
As shown in Table 5, only a small portion of dialogues can be fully understood, particularly for desire-related questions. This highlights that LLMs still lack the ability to reason about the dynamics of mental states in persuasive dialogues, resulting in a significant performance gap compared to humans.
@@ 5.3 How Well LLMs Track Mental States of Persuadees?
Qwen2.5-7B-Chat 19.12 31.81 56.95
Qwen2.5-7B-Chat + CoT 20.52 24.76 54.67
@@ 6 Conclusion
We introduce PersuasiveToM, a benchmark designed to evaluate the Machine Theory of Mind ability of LLMs in persuasive dialogues. The core of PersuasiveToM lies in the design of questions that probe the beliefs, desires, and intentions of both parties that have asymmetrical social status in real-life social interactions. Furthermore, we propose evaluating the ability to persuade based on understanding mental states. We conduct extensive experiments and analysis to evaluate the performance of LLMs on PersuasiveToM benchmark.
7 Limitations
@@
8 Societal and Ethical Considerations
@@
Appendix A Data Annotation
@@
Binary choice questions, including those related to the desires of the persuader and judgment questions, do not require additional choice generation. For belief-related questions for both parties, we adapt the tones from the PersuasionDaily dataset. We also create lists of attitudes—positive, neutral, and negative—and manually remove any items that have semantics too close to the ground truths. From each attitude list, we then randomly sample one word to generate the four choices.
@@
Appendix B Appendix for Experiments
B.1 Human Performance
@@
B.2 Prompts used for evaluation
Here we show the prompts used for vanilla zero-shot prompting and CoT prompting for generating answers for all the ToM Reasoning and ToM Application questions. We only need to fill in the content to "<>" for evaluating different questions. The vanilla zero-shot prompt is shown in Figure 9 and the CoT prompt is shown in Figure 10
@@ B.2 Prompts used for evaluation
B.3 Case study on Persuader’s intention
@@
Appendix C Details on Persuasive Principles

### v2

@@ 4.1 Baseline Models
We evaluate PersuasiveToM on eight frontier LLMs: Llama-3.1-8B-Instruct Dubey et al. (2024), Qwen-2.5-7B-Chat Yang et al. (2024), Gemma-2-9B-it Team et al. (2024), GLM4-9B-Chat GLM et al. (2024), Mixtral-8x7b-Instruct Jiang et al. (2024), and ChatGPT-series(GPT-4o-mini, GPT-4o-0806). By following the common practices Kim et al. (2023); Sabour et al. (2024), we test these models with two types of prompts: (1) vanilla zero-shot prompting directly asks LLMs to give a choice without any explanation; (2) CoT prompting method by following Kojima et al. (2022) and using the prompt “Let’s think step by step.” to elicit the reasoning process and extract the choices by string matching. The temperature for generating answers is set to 0.7111LLMs occasionally output with illegal format. We choose a low but nonzero temperature to resample the answers for these invalid generations., and we report results across three repetitions. To measure the specific performance gap between humans and the state-of-the-art machine on the PersuasiveToM, we employ three graduate students in computer science to complete the human evaluation task. To avoid the bias of LLMs toward a specific choice letter, we shuffle the choices to maintain a nearly uniform distribution of correct choices over the dataset. Prompts used for vanilla zero-shot prompting and CoT prompting are shown in Appendix C.2.
@@ 5.1 Main Results
Our results show that smaller models, such as Gemma-2-9B and Qwen-2.5-7B, can perform reasonably well in inferring the persuader’s desires, achieving an accuracy of over 96%, which is competitive with GPT-4o. This suggests that most LLMs can easily discern the desires of the persuader. However, when it comes to the desires of the persuadee, performance is relatively lower. Unlike the static desires of the persuader, the persuadee’s desires are dynamic, evolving from an initial state to a final state, often with neutral expressions in between. This lower performance highlights that inferring the dynamic desires of the persuadee remains a significant challenge for LLMs.
@@
On belief questions, larger models like GPT-4o perform much better than smaller models on reasoning about the beliefs of both parties. The performance difference between the reasoning persuader’s beliefs and the persuadee’s beliefs is subtle. This is because both parties’ beliefs dynamically change with each other’s speech during the conversation. The difficulty of reasoning persuader and the persuadee’s beliefs is similar.
@@
Strategy Prediction and Judgment.
@@
Both ToM reasoning and ToM application tasks indicate that CoT reasoning has not consistently improved performance, as observed in  Kim et al. (2023); Chen et al. (2024b), while it improves strategy prediction to some extent for most LLMs. CoT reasoning involves breaking down the mental states associated with each utterance and generating the mental states of previous utterances. This process can introduce intermediate mistakes that may mislead the overall reasoning about mental states.
@@
To obtain a baseline for human performance, we recruited participants to complete the questions. More details of human evaluation are shown in Appendix C.1. As shown in Table 5, our human participants outperformed LLMs on all tasks. In particular, although GPT-4o reaches close performance in humans, it still falls short of understanding and reasoning the complex dynamics such as the intention of persuaders and the desire of persuadees, which involves complex psychological changes, highlighting a significant gap in current LLMs and humans.
@@ 5.2 In-depth Analysis
Figure 3 summarizes the distribution of errors of desire questions happening in different stages of dialogue progress with and without CoT reasoning. The error distribution for persuader and persuadee is significantly different. At the beginning of the dialogue, LLMs may not accurately understand the persuader’s desire, but as the dialogue progresses, the persuader’s desire becomes relatively easy to identify. However, for the persuadee, the desire to reject at the early stage of the dialogue is relatively easy to recognize. As the persuasion proceeds, the persuadee may begin to contemplate and hesitate over the persuader’s proposal, leading to complex and nuanced psychological activities that make it difficult for the LLM to accurately judge the persuadee’s desire. As the dialogue approaches its end, the persuadee shows a tendency to agree, making the reasoning of desire easier. This suggests LLMs still fall short in ToM reasoning regarding desire shifts.
@@
Figure 4 and 5 summarize error types of belief questions for each model with and without CoT. We use Distil-BERT [^2] https://huggingface.co/distilbert/distilbert-base-uncased-finetuned-sst-2-english Sanh (2019) to discriminate whether the choice of LLMs has the same attitude polarity as the ground-truth. we found that LLMs make nearly balanced proportions of same- and opposite-polarity errors in both persuader and persuadee belief reasoning. This balance indicates that LLMs do not adopt a clear or consistent strategy when reasoning beliefs, often guessing without a coherent reasoning framework.
@@
Given the high error rate in intention questions related to the persuader, we conducted an analysis of the error types. Our findings reveal that most LLMs exhibit a bias toward predicting intentions characterized by making the other person feel accepted through concessions, promises, or benefits. Figure 6 illustrates the proportion of errors resulting from misclassifying intentions into this category. We hypothesize that this bias may stem from the pretraining phase, particularly with Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2017), which tends to prioritize safety and politeness. This may explain the models’ bias toward predicting intentions emphasizing benefits and concessions, even when misaligned with the dialogue context. We provide a case study in Appendix C.3.
Discussion.
LLMs struggle with BDI reasoning because these mental states are highly dynamic and context-dependent, making their accurate prediction across a dialogue challenging. To mitigate this issue, future work can explore incorporating external memory mechanisms to track changes in mental states over time. or focus on enhancing the LLM’s inherent understanding and reasoning capabilities regarding dynamic psychological states by constructing and training on larger, more richly annotated datasets that capture the nuanced evolution of beliefs, desires, and intentions in various interactive scenarios.
5.3 How Well LLMs Track the Mental States of Persuadees?
We evaluate whether LLMs can holistically track the persuadee’s mental states throughout the entire dialogue. To assess this, we measure whether the model maintains a coherent understanding across multiple turns, counting a dialogue as successful only if all related questions are answered correctly.
As shown in Table 6, only a small portion of the dialogues are fully understood, especially those involving desire reasoning. This indicates that LLMs still struggle to track dynamic mental states in persuasive settings, revealing a substantial gap from human performance.
@@ 5.3 How Well LLMs Track the Mental States of Persuadees?
Qwen2.5-7B-Instruct 19.12 31.81 56.95
Qwen2.5-7B-Instruct + CoT 20.52 24.76 54.67
@@ 6 Conclusion
This work proposes PersuasiveToM, a benchmark for evaluating LLMs’ ability to reason about complex psychological states and apply them in social decision-making. We conducted extensive experiments and analysis to evaluate the performance of LLMs on the PersuasiveToM benchmark.
Limitations
@@
Societal and Ethical Considerations
@@
Appendix A Additional Discussion
Dialogue Geneartion.
PersuasiveToM is built on multi-turn persuasive dialogues that require models to track changing beliefs, desires, and intentions of both the persuader and persuadee. In practice, LLMs often “get lost” in long conversations Yang et al. (2025); Laban et al. (2025). By evaluating how well LLMs can track evolving desires, beliefs, and intentions, PersuasiveToM provides insights into their capacity for generating contextually appropriate and coherence responses across multiple turns, as demonstrated in Section 5.
Personalized Dialogue.
PersuasiveToM assesses LLMs’ ability to reason about individual mental states, which is essential for tailoring responses to meet specific user needs and preferences Wang et al. (2024); Chen et al. (2024a). This understanding allows for more empathetic and effective interactions, enhancing user experience in applications such as customer service and mental health support Lee et al. (2022); Ma et al. (2023). Future work can develop external memory to store user’s mental states and retrieve form the memory to generate personalized responses Huang et al. (2024), or internalize them into the LLMs Li et al. (2023).
Appendix B Data Annotation
@@
Binary choice questions, including those related to the desires of the persuader and judgment questions, do not require additional choice generation. For belief-related questions for both parties, we adapt the tones from the DailyPersuasion dataset. We also create lists of attitudes—positive, neutral, and negative—and manually remove any items that have semantics too close to the ground truths. From each attitude list, we then randomly sample one word to generate the four choices.
@@
Appendix C Appendix for Experiments
C.1 Human Performance
@@
C.2 Prompts used for evaluation
Here we show the prompts used for vanilla zero-shot prompting and CoT prompting for generating answers for all the ToM Reasoning and ToM Application questions. We only need to fill in the content to "<>" for evaluating different questions. The vanilla zero-shot prompt is shown in Figure 9, and the CoT prompt is shown in Figure 10
@@ C.2 Prompts used for evaluation
C.3 Case study on Persuader’s intention
@@
Appendix D Details on Persuasive Principles